Find the interval(s) on which f (x) = 2x2 – 3x2 – 12x + 1 is concave up.
Answer #1

f(x) = 2x² – 3x² - 12a + 1. f(x) = 6x² - 6x - 12 f(x) = 12x– 6 putting f(x) = 0 → 12x-6 = 0 checking the value of tcx) on